OpenStudy (anonymous):

P^2-4p-32/p+4 @ghazi –p + 8; p ≠ –4 p – 8; p ≠ –4 –p – 8; p ≠ –4 p + 8; p ≠ –4

OpenStudy (ghazi):

well \[P^2-4p-32=(p-8)(p+4)\]

OpenStudy (ghazi):

\[\frac{ (p-8)(p+4) }{ (p+4) }=(p-8)\] p=/ 4

OpenStudy (anonymous):

confused

OpenStudy (ghazi):

@Gungirl4ever1994 clear?

OpenStudy (anonymous):

i no =/4 bc thats at the end of em all

OpenStudy (ghazi):

sorry \[p \ne 4\]

OpenStudy (ghazi):

b

OpenStudy (anonymous):

bubba help on more?

OpenStudy (ghazi):

keep posting :)

OpenStudy (anonymous):

???

OpenStudy (ghazi):

\[\frac{ q^2+11q+24 }{ q^2-5q-24 } =\frac{ (q+8)(q+3) }{ (q-8)(q+3) }=\frac{ (q+8) }{ (q-8) }\]

OpenStudy (ghazi):

C

OpenStudy (ghazi):

you can do this one :) you need to factorize

OpenStudy (anonymous):

plzzz

OpenStudy (ghazi):

yes you can do it :D z^2+3z+2=(z+2)(z+1)

OpenStudy (ghazi):

A

OpenStudy (anonymous):

thanks talk to you in a bit :)